OpenStudy (anonymous):

simplify: tan x/(secx-1)

OpenStudy (anonymous):

Multiply by \(\dfrac{\sec x+1}{\sec x+1}\) and try reducing that.

OpenStudy (anonymous):

I basically just get the opposite: secx+1/tanx

OpenStudy (anonymous):

Well, the instructions are kinda vague. Did you learn any half-angle identities recently? This might be an application of that.

OpenStudy (anonymous):

We haven't learned those - this is part of a summer review packet and it doesn't say at all how she wants us to do it. Normally it's just using trig identities

OpenStudy (loser66):

that's good, from then, you have sec / tan + 1/tan = 1/sin + cot = csc + cot

OpenStudy (anonymous):

oh ok what @Loser66 makes sense. Thanks!

OpenStudy (loser66):

yw

OpenStudy (anonymous):

It's also equal to \(\tan\left(\dfrac{x}{2}\right)\), but again, not sure if that's what the packet's looking for as an answer.

OpenStudy (anonymous):

Thanks, but I think that the cscx+cotx was more what it was looking for